A spring is such that a 2 lb weight stretches it by 6 in. There is a damping force present, with



the magnitude equal to 0.75|𝑣|. An impressed force of F(t) = 2 sin 8t is acting on the spring. If



the 2-lb weight is released from a point 3 in below the equilibrium point, describe the motion

1
Expert's answer
2021-12-13T14:54:02-0500

Momentum after release of steam P.F.f 

Pf = mv + m'v'       (1)

Where mass of empty launch steam m=7500 kg

m' is mass of satellite , m'=250 kg

v is final velocity of launch steam after releasing satellite, v=900 m/s

v' is final velocity of the satellite after release

Putting all the values in equation (1), we get

Pf =(7500 kg)(900 m/s) + (250 kg)v'

Pf = 6.750×10kg•m/s + (250 kg)v'


Now applying conservation of momentum

Momentum before release of satellite = momentum after release of satellite

Pi = Pf

7.750×106 kg•m/s = 6.750×10kg•m/s + (250 kg)v'

(7.750×106 - 6.750×10)kg•m/s= (250 kg)v'

106 m/s = 250v'

v' = (106/250) m/s

v' = 4000 m/s